For a holiday party, Laura is making a seven-layer torte, which consists of seven 8-inch diameter cake layers with a cream topping on each layer. She knows that one recipe of cream will cover 55 square inches of cake. How many batches of the topping must she make to complete her torte? (Use п = 3.14 as an approximate value)

I do not understand what they are asking?

Answer by rfer(16322)

Find the area of each cake, multiply by seven, the divide by 55 for total batches.
A=pir^2
A=3.14*4^2
A=3.14*16
A=50.24 sq in each cake layer
-------------------
50.24*7=351.68 sq in
------------
351.68/55=6.394 batches
she would need 7 batches of cream